A production process consists of a three-step operation. The scrap rate is 11 percent for the first step and 8 percent for the other two steps.

a.If the desired daily output is 495 units, how many units must be started to allow for loss due to scrap? (Do not round intermediate calculations. Round up your final answer to the next whole number.)

Number of units           

b.If the scrap rate for each step could be cut in half at every operation, how many units would this save in terms of the scrap allowance? (Do not round intermediate calculations. Round up your final answer to the next whole number.)

Number of units           

c.If the scrap represents a cost of $10 per unit, how much is it costing the company per day for the original scrap rate (i.e. the Part a scrap rate)? (Round your final answer to the nearest whole number. Omit the "$" sign in your response.)

Cost           $
Answer #1

Answer:

  • Scrap rate for whole process = 1 - [ (1- rate of 1st) * (1- rate of 2nd)* (1-rate of 3) ]
  • Scrap rate for whole process = 1 - [ (1-0.11)*(1-0.08)*(1-0.08)] = 0.246704

a)

  • the % of correct units = 1 - scrap rate = 75.32 %
  • Daily Demand is 495
  • So 495 = 75.32% of the production
  • Total production = 495 / 0.7532 =657.19
  • So total nunits to start with = 658

B)

  • New overall scrap rate = 1 - [ (1-0.055)*(1-0.04)*(1-0.04)] = 0.1291
  • the % of correct units = 1 - scrap rate = 87.09%
  • Daily Demand is 495
  • So 495 = 87.09% of the production
  • Total production = 495 / 0.8709 =568.369
  • Total units saved = total starting unit previously - total starting now = 89
  • Answer = 89

c)

  • total units scraped = total start units - total end units = 657.19 - 495 = 162.19
  • Total cost = 162.19*10 = 1621.9 = 1622 (Answer)